OpenStudy (anonymous):

(sqrt 3x-2) +3=4x ? Can someone help me with this question ?

jhonyy9 (jhonyy9):

like a first add to the both sides minus 3 and in the second step square both sides to elimine the squarroot ok ?

jhonyy9 (jhonyy9):

sqrt(3x-2) +3 = 4x this is correct wrote ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so can you calculi how i have wrote it above step by step ?

jhonyy9 (jhonyy9):

comen this is easy

jhonyy9 (jhonyy9):

courage

jhonyy9 (jhonyy9):

than you add to the both sides minus 3 what will get ?

jhonyy9 (jhonyy9):

courage i will help you but you need to cooperate anything

jhonyy9 (jhonyy9):

sqrt(3x-2) +3 = 4x / + (-3) sqrt(3x-2) +3 -3 = 4x -3 do you can continue it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you sqrt both sides by 2 ?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

you have got the result ?

OpenStudy (anonymous):

I got stuck after sqrting it by 2

jhonyy9 (jhonyy9):

why ? please write the calculi here

OpenStudy (anonymous):

(sqrt3x-2)^2 = (4x-3)^2 3x-2 =4x-3 ?

jhonyy9 (jhonyy9):

the left side is right but the right side not is you dont squared the (4x-3) ,why ?

OpenStudy (anonymous):

how do i do that?

jhonyy9 (jhonyy9):

(4x-3)^2 =(4x-3)(4x-3) yes ?

OpenStudy (anonymous):

ohh okay i get it now

jhonyy9 (jhonyy9):

can you get the right result now sure ?

OpenStudy (anonymous):

let me try hahah

OpenStudy (anonymous):

do i add the right side together ?

jhonyy9 (jhonyy9):

no there is (4x-3) time (4x-3)

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

okay what ?

jhonyy9 (jhonyy9):

you dont like writing the calculi here ?

OpenStudy (anonymous):

3x-2=16x^2-6x+9

OpenStudy (anonymous):

is that correct ?

jhonyy9 (jhonyy9):

not is correct the right side

OpenStudy (anonymous):

it's supposed to be 8x^2 ?

jhonyy9 (jhonyy9):

so (4x-3)*(4x-3)= 4x*4x -3*4x -3*4x +(-3)*(-3) = ?

jhonyy9 (jhonyy9):

do you see why your calculi not is right ?

jhonyy9 (jhonyy9):

so 4x*4x = 16x^2 is correct so than continue -3*4x =-12x and -3*4x=-12x so -12x -12x = -24x yes ? and (-3)*(-3) = +9 than now do you see where not is correct your calculi ?

OpenStudy (anonymous):

ohhhh I understand now

jhonyy9 (jhonyy9):

than do you can finish it now easy ?

OpenStudy (anonymous):

0=16x^2-9x-7 ?

jhonyy9 (jhonyy9):

than 3x-2=16x^2 -24x +9 yes ?

jhonyy9 (jhonyy9):

how you get -9x and -7 ?

OpenStudy (anonymous):

16x^2 - 24x + 9 -3x-2

jhonyy9 (jhonyy9):

why -2 ?

OpenStudy (anonymous):

so the other side can equal to 0

jhonyy9 (jhonyy9):

but the -2 not is right

jhonyy9 (jhonyy9):

because on the left side there are 3x-2 so than thiis on the right side will be -3x+2 yes ?

OpenStudy (anonymous):

ahhhh okay

OpenStudy (anonymous):

so instead of 7 it will be 11 ?

jhonyy9 (jhonyy9):

yes and the term of x how will be ?

jhonyy9 (jhonyy9):

first please write it correct the quadratic what you need to get

jhonyy9 (jhonyy9):

so 16x^2 and how you continue it ?

OpenStudy (anonymous):

I don't know haha

jhonyy9 (jhonyy9):

so there are -24x-3x yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

than how will be the resulted quadratic correct ?

jhonyy9 (jhonyy9):

you dont like continue it ?

OpenStudy (anonymous):

sorry I was going over my problem

OpenStudy (anonymous):

0=16^2-21x+11

jhonyy9 (jhonyy9):

ok so than how will be ?

jhonyy9 (jhonyy9):

there are -24x-3x

OpenStudy (anonymous):

which is 21x

jhonyy9 (jhonyy9):

-24-3 why is it 21 ?

OpenStudy (anonymous):

ohh sorry it's 27 right ? because of the - sign

OpenStudy (anonymous):

I'm sorry I'm stupid in math.

jhonyy9 (jhonyy9):

so than how will be the quadratic correct ? and you need to solve it ?

OpenStudy (anonymous):

(4x-3)(x-1)

jhonyy9 (jhonyy9):

no begin it with 16x^2

OpenStudy (anonymous):

16x^2 -27x+11 is this what your talking about ?

jhonyy9 (jhonyy9):

yes than this quadratic you need to solve it ,yes ?

OpenStudy (anonymous):

that would be the answer ?

jhonyy9 (jhonyy9):

i dont know do you know how to yolve a quadratic ?

OpenStudy (anonymous):

Yes

jhonyy9 (jhonyy9):

so than you need to solve it how ? i think you need to us the discriminant method

jhonyy9 (jhonyy9):

because you not can factoriz it ,yes ?

jhonyy9 (jhonyy9):

yes this is right but in case of your exercise how will be first you need to calculi the roots of quadratic

OpenStudy (anonymous):

I don't know ?

jhonyy9 (jhonyy9):

us the discriminant method do you know it ? D=b^2 -4ac

jhonyy9 (jhonyy9):

and x1,2 = ?

OpenStudy (anonymous):

No I haven't learn it

jhonyy9 (jhonyy9):

so than i think that you dont write correct your exersice above how is it in your book ?

OpenStudy (anonymous):

I checked it's correct

OpenStudy (anonymous):

I was saying that I haven't learned about the discriminant method

jhonyy9 (jhonyy9):

but this qudratic you can solve it using this method

OpenStudy (anonymous):

ohh but I don't know how to use that method.

jhonyy9 (jhonyy9):

sorry but than i think that you need to learn it for you can solve this exersice

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

good luck bye